This question is not from a text book. Therefore, I don't have anything to provide.
Please solve this in a step-by-step format:
The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.

Found 2 solutions by checkley71, Nate:
Answer by checkley71(8403)

B=BASE THEN 2(4B-8)+B=29 OR 8B-16+B=29 OR 9B=45 OR B=45/9 OR B=5
PROOF 2(20-8)+5=29 OR 2*12+5=29 OR 24+5=29 OR 29=29

Answer by Nate(3500)

Base's side length: b
Leg's side length: 4b - 8
side + side + side = perimeter
b + (4b - 8) + (4b - 8) = 29
9b - 16 = 29
9b = 45
b = 5
Leg's side length:
4b - 8
4(5) - 8
20 - 8
12
Side's lengths: 5,12,12